# Zhixuan Chu1∗♠, Yan Wang1∗, Longfei Li1, Zhibo Wang2, Zhan Qin2, Kui Ren2

### v1

@@ 5.4 Main Results
5.5 Ablation Study
5.5.1 Different Components
Table 2 presents the ablation studies conducted to evaluate the impact of different components in the LLMGuardaril framework. The experiments are performed on two model variants, Llama2-7b and Llama2-13b, using three benchmark datasets: ToxiGen, BOLD, and AdvBench. The ablation studies focus on three key components of LLMGuardaril: Causal Debias Loss: This loss term is designed to debias the influence of the semantic direction representation on the direction representations during the adversarial learning process. By minimizing this loss, LLMGuardaril aims to mitigate the biases from the steering process and obtain unbiased steering representations. Prediction Loss: The prediction loss ensures that the original semantic information remains unchanged during the debiasing process. It measures the cross-entropy between the original output generated by the LLM using the original representations and the output generated using the debiased representations. Minimizing this loss helps maintain the model’s general knowledge and capabilities while applying the steering. Intervened Layer Selection: LLMGuardaril employs a systematic approach to identify the most effective layers for intervention based on probing accuracy. The selected layers are then used in the Debias LoRA Block to obtain the debiased representations and calculate the steering representation. Without this intervened layer selection module, we directly select the middle layers, for example, 6 layers for Llama2-7B (total 32 layers) and 10 layers for Llama2-13B (total 40 layers).
The ablation studies are conducted by removing each component individually and evaluating the model’s performance. The results are compared to the base model and the complete LLMGuardaril framework. Overall, the ablation studies demonstrate the significance of each component in LLMGuardaril. The Causal Debias Loss is crucial for obtaining unbiased steering representations and effectively mitigating undesired attributes. The Prediction Loss plays a vital role in maintaining the model’s general knowledge and ensuring the quality of the generated outputs. The Intervened Layer Selection contributes to the overall performance by identifying the most effective layers for intervention. The complete LLMGuardaril framework, incorporating all these components, achieves the best results across the benchmark datasets, validating the effectiveness of our proposed approach.
Table 3: Ablation studies of our LLMGuardaril for different output control operation choices. $\uparrow$ means higher is better and $\downarrow$ means lower is better.
@@ 5.5.1 Different Components
5.5.2 Different Output Control Operations
Table 4: Scaling Law studies of our LLMGuardaril model for different-sized LLMs. $\uparrow$ means higher is better and $\downarrow$ means lower is better.
@@ 5.5.2 Different Output Control Operations
Vicuna-7b LLMGuardaril (ours) 85.59 14.02 59.55 0.738 86.76 12.90
@@
Vicuna-13b LLMGuardaril (ours) 86.60 13.33 60.35 0.740 87.22 12.60
@@
Vicuna-33b LLMGuardaril (ours) 87.45 12.34 61.22 0.749 87.96 12.00
5.6 The Study of Scaling Law
Table 4 presents the scaling law study conducted to evaluate the performance of LLMGuardaril across different model sizes within the Vicuna family. The experiments are performed on three benchmark datasets: ToxiGen, BOLD, and AdvBench. The base model performance is compared to LLMGuardaril for each model size: Vicuna-7b, Vicuna-13b, and Vicuna-33b. The scaling law study aims to investigate the effect of increasing model size on the performance of LLMGuardaril in steering the model’s output towards desired attributes. It explores the benefits of LLMGuardaril scale with the model size and if larger models exhibit better alignment with the target attributes.
Overall, the scaling law study demonstrates the robustness and scalability of LLMGuardaril across different model sizes within the Vicuna family. It shows the benefits of LLMGuardaril in steering the model’s output towards desired attributes scale with the model size, providing insights into the relationship between model capacity and steering performance. The results suggest that LLMGuardaril can be effectively applied to even larger models to achieve better alignment with target attributes while mitigating undesired behaviors.
Table 5: Out-of-domain (OOD) experiments of our LLMGuardaril model on the ToxiGen dataset. $\uparrow$ means higher is better and $\downarrow$ means lower is better.
@@ 5.6 The Study of Scaling Law
LLMGuardaril (ours) 89.54 10.41 80.87 18.17 100 0 73.17 25.00 100 0
5.7 OOD Experiments of LLMGuardaril
@@
Table 5 presents the results of out-of-domain (OOD) experiments for the LLMGuardaril model and several baseline methods on the ToxiGen dataset. The experiments aim to assess the model’s ability to generalize the steering representations learned from one demographic group (women) to other demographic groups (Black, Muslim, Native American, Latino, and Jewish).
@@
LLMGuardaril demonstrates strong generalization capability by effectively reducing toxicity and increasing refusal rates for all target demographic groups, even though the steering representations were learned using examples from the women group. This suggests that the learned steering representations capture general patterns of toxicity and harmfulness that can be applied to different demographic contexts. In addition, LLMGuardaril consistently outperforms the baseline methods (Base, LAT-Reading, LORRA, and Mean-Centring) across all target demographic groups, achieving higher refusal rates and lower toxicity percentages. This highlights the effectiveness of the causal analysis and adversarial learning techniques employed by LLMGuardaril in obtaining unbiased steering representations that generalize well to unseen demographic groups.
Figure 6: Sensitivity analysis of $\alpha$.
Figure 7: Sensitivity analysis of $\beta$.
5.8 Sensitivity Analysis of LLMGuardaril
@@
Figure 6 and 7 present the sensitivity analysis of the hyperparameter $\alpha$ and $\beta$ on the ToxiGen and AdvBench datasets. The hyperparameter $\alpha$ controls the balance between the prediction reconstruction loss and the debias loss in the overall loss function. The findings suggest that the prediction reconstruction loss is a vital element in the training process and carries more weight compared to the debias loss in terms of maintaining the model’s effectiveness. The hyperparameter $\beta$ represents the intervention strength when controlling the model output using the steering representations. Based on the trends observed in the sensitivity analyses, an optimal value for $\beta$ appears to be around 2. This value achieves a good balance between refusing harmful prompts and minimizing toxic outputs while maintaining the model’s performance. The optimal ranges identified for $\alpha$ and $\beta$ can guide the selection of appropriate values for these hyperparameters in practical applications of the LLMGuardaril model.
6 Discussion
In this paper, we introduced LLMGuardaril, a novel framework for obtaining unbiased steering representations in large language models (LLMs) by incorporating causal analysis and adversarial learning techniques. Our approach aims to mitigate the influence of semantic biases on the steering process and provide explainable insights into the generated output. In this section, we discuss the potential social impact of our work, address the limitations of our framework, and outline future research directions.
6.1 Potential Social Impact
The development of safe and reliable LLMs has significant social implications. As these models become increasingly prevalent in various domains, such as content generation, dialogue systems, and decision support, it is crucial to ensure that their outputs align with desired attributes and do not perpetuate harmful biases. Our proposed LLMGuardaril framework contributes to this goal by enabling the steering of LLMs toward desired directions while mitigating the influence of unwanted biases.
The explainability component of LLMGuardaril has the potential to foster trust and transparency in the use of LLMs. By providing insights into the alignment between the generated output and the desired direction, our framework allows users to understand the specific aspects of the output that contribute to the desired attributes. This explainability feature can be particularly valuable in sensitive domains, such as healthcare, finance, and legal systems, where the ability to interpret and validate the model’s outputs is essential.
Furthermore, the unbiased steering representations obtained through LLMGuardaril can help mitigate the propagation of societal biases and stereotypes in the generated text. By disentangling the influence of semantic biases from the steering process, our framework can contribute to the development of more equitable and inclusive language models. This has the potential to promote fairness and reduce discrimination in applications that rely on LLMs for decision-making or content generation.
However, it is important to acknowledge that the impact of LLMs on society is multifaceted and requires ongoing research and ethical considerations. While frameworks like LLMGuardaril can help in steering LLMs towards desired attributes, it is crucial to ensure that the desired directions themselves are carefully defined and aligned with societal values and norms. The responsible development and deployment of LLMs necessitate a collaborative effort among researchers, policymakers, and stakeholders to address the broader ethical implications.
6.2 Limitations
While LLMGuardaril demonstrates promising results in obtaining unbiased steering representations and providing explainable insights, there are certain limitations to our framework that should be acknowledged.
First, the effectiveness of LLMGuardaril relies on the availability and quality of the steering prompts used to represent the desired and undesired attributes. Constructing appropriate prompts that accurately capture the intended steering direction can be challenging and may require domain expertise. The performance of our framework may be sensitive to the choice of prompts, and poorly designed prompts can lead to suboptimal steering results.
Second, the adversarial learning approach employed in LLMGuardaril involves a training process that can be computationally intensive, especially for large-scale LLMs. The training time and resource requirements may pose practical constraints on the applicability of our framework in certain scenarios. Further research is needed to optimize the training process and explore more efficient techniques for obtaining unbiased steering representations.
Third, while LLMGuardaril aims to mitigate the influence of semantic biases on the steering process, it may not completely eliminate all forms of bias. The framework relies on the assumption that adversarial learning can effectively disentangle the semantic biases from the steering representations. However, there may be residual biases or subtle interactions that are not fully captured by our approach. Continuous evaluation and refinement of the framework are necessary to address potential limitations and improve its robustness.
6.3 Future Work
The proposed LLMGuardaril framework opens up several avenues for future research in the field of language model steering and explainable AI. Here, we outline some potential directions for further exploration.
First, extending LLMGuardaril to handle multiple steering attributes simultaneously is an important direction. In real-world scenarios, it may be desirable to steer LLMs towards multiple desired attributes or concepts concurrently. Investigating techniques to incorporate multiple steering representations and develop appropriate weighting schemes to balance their influence on the generated output is a valuable research direction.
Second, exploring the integration of LLMGuardaril with other approaches for controlling the output of LLMs, such as fine-tuning or prompt engineering, can lead to more comprehensive and effective steering strategies. Combining the strengths of different approaches may yield improved performance and greater flexibility in guiding the model’s output towards desired attributes.
Third, further research on the explainability component of LLMGuardaril can focus on developing more advanced techniques for analyzing the alignment between the generated output and the steering representations. Investigating methods to provide more fine-grained explanations, such as identifying specific phrases or patterns that contribute to the desired or undesired attributes, can enhance the interpretability of the steering process.
Fourth, conducting user studies to evaluate the usability and effectiveness of LLMGuardaril in real-world applications is crucial. Engaging with domain experts and end-users to gather feedback on the explainability and control aspects of the framework can provide valuable insights for improvement and adaptation to specific use cases.
Finally, addressing the broader ethical implications of language model steering and explainable AI is an important research direction. Collaborating with researchers from diverse fields, including ethics, social sciences, and law, can help in developing guidelines and best practices for the responsible development and deployment of LLMs. Engaging in multidisciplinary research efforts can contribute to the development of AI systems that are not only effective but also aligned with societal values and norms.
In conclusion, LLMGuardaril presents a novel framework for obtaining unbiased steering representations in LLMs by incorporating causal analysis and adversarial learning techniques. Our approach aims to mitigate the influence of semantic biases on the steering process and provide explainable insights into the generated output. While the framework demonstrates promising results, there are limitations and opportunities for future research. The potential social impact of our work highlights the importance of developing safe and reliable LLMs that align with desired attributes and promote fairness and inclusivity. Ongoing research and collaborative efforts are necessary to address the broader ethical implications and ensure the responsible development and deployment of language models in real-world applications.
7 Conclusion

### v2

@@ 5.1. Baselines
Table 3. Ablation studies of our LLMGuardrail for different output control operation choices. $\uparrow$ means higher is better and $\downarrow$ means lower is better.
@@
Figure 5. The examples of the prefix steering prompt sets, and the original and intervened outputs by our LLMGuardrail with explainable shading.
5.2. Benchmarks and Evaluation Metrics
To assess the effectiveness of LLMGuardrail in steering language models toward desired attributes, we conduct experiments on several public datasets that focus on different aspects of content safety and bias (Wang and Shu, 2023). These datasets are carefully selected to cover a wide range of challenging scenarios and provide a comprehensive evaluation of our proposed framework.
[…]
Toxicity. To evaluate the effectiveness of LLMGuardrail in mitigating toxicity, we employ the ToxiGen dataset (Hartvigsen et al., 2022), which contains implicitly toxic and benign sentences mentioning 13 minority groups. We use a revised version of the dataset (Hosseini et al., 2023) that reduces noise by filtering out prompts with disagreement among annotators regarding the target demographic group. The main metric is the percentage of toxic generations, determined using HateBERT, a fine-tuned BERT model provided by the dataset. Furthermore, we report the percentage of refusal responses, identified by the presence of specific signal keywords, to assess the model’s ability to avoid engaging with potentially harmful prompts.
Bias. The BOLD benchmark (Dhamala et al., 2021) is employed to evaluate bias in generated responses. This large-scale dataset comprises 23,679 English Wikipedia prompts spanning five domains: race, gender, religion, political ideology, and profession. To manage experiment costs, we sample 120 prompts from each domain. The VADER sentiment score (Hutto, 2022) serves as the primary metric, quantifying the sentiment directed toward the population mentioned in each prompt. VADER generates a sentiment score between -1 and 1, with 0 indicating a neutral sentiment. While the goal is to identify imbalances in sentiment across different groups, for conciseness, we report the mean sentiment score over the entire dataset as our main metric. Additionally, we provide the percentage of refusal responses to assess the model’s ability to avoid biased or discriminatory language.
Harmfulness. To evaluate LLMGuardrail’s performance in mitigating harmful content, we utilize the AdvBench dataset (Zou et al., 2023b), which contains 500 harmful behaviors and instructions reflecting toxic, discriminatory, or cybercrime-related actions. The primary metric is the percentage of refusal responses, identified using the same key phrases for refusal as in the original dataset. Additionally, we employ HateBERT, a fine-tuned BERT model, to classify the toxicity of generated responses (excluding refusals). The percentage of toxic generations serves as an additional metric to assess the model’s ability to avoid producing harmful content.
By conducting experiments on these diverse benchmarks, we aim to provide a comprehensive evaluation of LLMGuardrail’s effectiveness in steering language models toward desired attributes while mitigating undesirable behaviors. The selected datasets cover a wide range of content safety and bias challenges, enabling us to assess the framework’s performance in promoting truthfulness, reducing toxicity and bias, and avoiding harmful content generation.
5.3. Experiment Settings
Prompt Design. To thoroughly assess the efficacy of LLMGuardrail, we employ a prompt structure that combines a prefix steering prompt, which can be either positive or negative, with an identical semantic prompt. This approach enables us to isolate the influence of the steering prompt on the model’s generated output while maintaining a consistent semantic context across experiments. As depicted in Figure 5, we meticulously develop the prefix steering prompt sets tailored to each benchmark dataset: TruthfulQA, BOLD, ToxiGen, and AdvBench. These prompts are strategically designed to steer the model toward generating content that aligns with the desired attributes or concepts specific to each dataset, such as truthfulness, lack of bias, non-toxicity, and avoidance of harmful content.
Model Selection. We evaluate the guardrail performance of LLMGuardrail using two prominent families of instruction-tuned large language models: Llama2 (Touvron et al., 2023) and Vicuna-V1.5 (Chiang et al., 2023). The choice of these models is motivated by their widespread adoption and exceptional performance. To strike a balance between computational feasibility and comprehensive evaluation, our primary experiments focus on the 7B and 13B variants within each model family. These model sizes provide a reasonable trade-off between performance and resource requirements, enabling us to conduct extensive experiments while managing computational costs effectively. To gain a deeper understanding of LLMGuardrail’s scalability and its potential to generalize across different model sizes, we expand our experiments to encompass the entire spectrum of model variants within the Vicuna family, including Vicuna-33B.
Table 4. Scaling Law studies of our LLMGuardrail model for different-sized LLMs. $\uparrow$ means higher is better and $\downarrow$ means lower is better.
@@ 5.3. Experiment Settings
Vicuna-7b LLMGuardrail (ours) 85.59 14.02 59.55 0.738 86.76 12.90
@@
Vicuna-13b LLMGuardrail (ours) 86.60 13.33 60.35 0.740 87.22 12.60
@@
Vicuna-33b LLMGuardrail (ours) 87.45 12.34 61.22 0.749 87.96 12.00
Table 5. Out-of-domain (OOD) experiments of our LLMGuardrail model on the ToxiGen dataset. $\uparrow$ means higher is better and $\downarrow$ means lower is better.
@@
LLMGuardrail (ours) 89.54 10.41 80.87 18.17 100 0 73.17 25.00 100 0
Table 6. Performance comparison with aligned LLMs by SFT, DPO, or PPO on four benchmark datasets.
BaseModel Method TruthfulQA ToxiGen BOLD AdvBench
Base 34.08 88.32 30.10 54.00 44.71 35.00 0.438 80.58 19.04
SFT 48.32 71.09 34.35 56.12 42.18 39.44 0.451 81.42 18.51
Vicuna-7b DPO 50.42 80.46 40.57 80.47 18.55 56.81 0.626 85.31 14.65
PPO 52.33 81.21 42.50 84.90 14.05 59.49 0.709 86.55 13.42
LLMGuardrail (ours) 44.74 95.63 42.78 85.59 14.02 59.55 0.738 86.76 12.90
Base 34.75 89.52 31.11 54.71 43.57 0.45 0.746 65.58 34.42
SFT 47.91 74.03 35.47 58.19 40.79 2.27 0.760 67.89 32.08
Llama2-7b DPO 49.78 76.60 38.13 81.77 18.02 6.34 0.882 79.05 20.92
PPO 51.62 77.59 40.05 86.24 13.52 7.85 0.890 81.60 18.36
LLMGuardrail (ours) 42.31 95.60 40.45 86.29 13.01 8.00 0.895 80.85 19.15
Base 45.33 90.80 41.15 57.57 42.43 3.57 0.863 68.46 30.77
SFT 47.92 90.04 43.15 60.38 39.00 4.02 0.856 68.11 31.70
Llama2-13b DPO 54.70 84.20 46.06 87.26 12.45 7.25 0.878 79.09 20.15
PPO 55.79 84.94 47.39 90.16 9.62 8.70 0.900 80.42 19.46
LLMGuardrail (ours) 48.22 96.77 46.67 88.85 10.15 8.64 0.899 80.96 19.04
Table 7. Out-of-domain (OOD) experiments, performance comparison with aligned LLMs on ToxiGen dataset.
Base 64.72 35.20 53.48 26.11 78.24 21.76 29.13 69.66 80.61 19.17
SFT 73.46 25.15 70.16 29.58 88.35 11.26 35.24 64.22 85.40 14.44
Llama2-7b DPO 84.47 14.80 73.98 25.87 93.96 5.28 53.90 45.92 95.37 4.49
PPO 88.93 10.36 74.25 24.80 97.28 2.60 55.95 43.55 95.45 4.30
LLMGuardrail (ours) 90.73 8.13 75.21 23.15 100 0 56.20 42.62 100 0
Base 68.55 31.45 55.37 44.63 84.49 15.51 35.44 61.50 81.31 17.66
SFT 75.72 24.21 70.62 28.85 89.96 10.00 38.59 61.20 87.90 11.87
Llama2-13b DPO 84.30 15.45 78.42 21.63 95.42 4.31 73.05 26.90 97.49 2.50
PPO 88.69 11.15 79.04 20.80 98.80 0.76 72.59 27.10 98.77 1.13
LLMGuardrail (ours) 89.54 10.41 80.87 18.17 100 0 73.17 25.00 100 0
5.4. Main Results
Table 1 presents a comprehensive performance comparison of LLMGuardrail against various baseline methods across four benchmark datasets: TruthfulQA, ToxiGen, BOLD, and AdvBench. These datasets evaluate the models’ ability to align with desired attributes such as truthfulness, non-toxicity, lack of bias, and avoidance of harmful content. Across all datasets and model variants (Vicuna-7b, Llama2-7b, and Llama2-13b), LLMGuardrail consistently outperforms the baseline methods. The results highlight LLMGuardrail’s state-of-the-art performance in steering the language models toward the desired attributes. It successfully mitigates undesirable behaviors and aligns the generated outputs with the target objectives. The superior performance of LLMGuardrail compared to the baseline methods highlights the significance of our proposed framework and its potential for real-world applications. As depicted in Figure 5, we provide examples comparing the original LLM output without any guardrails to the intervened output generated by our LLMGuardrail. The original and intervened outputs use shading to highlight tokens based on their degree of alignment with the desired direction, as measured by the similarity scores between the token representations and the steering representation.
5.5. Ablation Study
5.5.1. Different Components
Table 2 presents the ablation studies conducted to evaluate the impact of different components in the LLMGuardrail framework. The experiments are performed on two model variants, Llama2-7b, and Llama2-13b, using three benchmark datasets: ToxiGen, BOLD, and AdvBench. The ablation studies focus on three key components of LLMGuardrail: Causal Debias Loss: This loss term is designed to debias the influence of the semantic direction representation on the direction representations during the adversarial learning process. By minimizing this loss, LLMGuardrail aims to mitigate the biases from the steering process and obtain unbiased steering representations. Prediction Loss: The prediction loss ensures that the original semantic information remains unchanged during the debiasing process. It measures the cross-entropy between the original output generated by the LLM using the original representations and the output generated using the debiased representations. Minimizing this loss helps maintain the model’s general knowledge and capabilities while applying the steering. Intervened Layer Selection: LLMGuardrail employs a systematic approach to identify the most effective layers for intervention based on probing accuracy. The selected layers are then used in the Debias LoRA Block to obtain the debiased representations and calculate the steering representation. Without this intervened layer selection module, we directly select the middle layers, for example, 6 layers for Llama2-7B (total 32 layers) and 10 layers for Llama2-13B (total 40 layers).
The ablation studies are conducted by removing each component individually and evaluating the model’s performance. The results are compared to the base model and the complete LLMGuardrail framework. Overall, the ablation studies demonstrate the significance of each component in LLMGuardrail. The Causal Debias Loss is crucial for obtaining unbiased steering representations and effectively mitigating undesired attributes. The Prediction Loss plays a vital role in maintaining the model’s general knowledge and ensuring the quality of the generated outputs. The Intervened Layer Selection contributes to the overall performance by identifying the most effective layers for intervention. The complete LLMGuardrail, incorporating all these components, achieves the best results across the benchmark datasets, validating the effectiveness of our proposed approach.
Figure 6. Sensitivity analysis of $\alpha$.
Figure 7. Sensitivity analysis of $\beta$.
5.5.2. Different Output Control Operations
Figure 8. Visualization of t-SNE on the AdvBench subset dataset. Red indicates representations of prompts identified as unsafe, dark blue represents representations of prompts identified as safe, light blue denotes prompt representations generated by jailbreak attacks, and yellow indicates representations obtained by calibrating jailbreak attack prompts using LLMGuardrail.The visualization demonstrates how LLMGuardrail effectively mitigates jailbreak attacks. Originally, the jailbreak attack prompts (light blue) are positioned near the boundary between safe and unsafe prompts, potentially causing the model to misclassify them as safe. After applying LLMGuardrail, these prompts are shifted toward the unsafe region (yellow), enabling the model to correctly identify them as potentially harmful. This recalibration allows the model to properly defend against or refuse to answer unsafe queries, thereby improving its overall safety and reliability.
5.6. The Study of Scaling Law
Table 4 presents the scaling law study conducted to evaluate the performance of LLMGuardrail across different model sizes within the Vicuna family. The experiments are performed on three benchmark datasets: ToxiGen, BOLD, and AdvBench. The base model performance is compared to LLMGuardrail for each model size: Vicuna-7b, Vicuna-13b, and Vicuna-33b. The scaling law study aims to investigate the effect of increasing model size on the performance of LLMGuardrail in steering the model’s output toward desired attributes. It explores the benefits of the LLMGuardrail scale with the model size and if larger models exhibit better alignment with the target attributes. Overall, the scaling law study demonstrates the robustness and scalability of LLMGuardrail across different model sizes. It shows the benefits of LLMGuardrail in steering the model’s output toward desired attributes scale with the model size, providing insights into the relationship between model capacity and steering performance. The results suggest that LLMGuardrail can be effectively applied to even larger models to achieve better alignment with target attributes while mitigating undesired behaviors.
5.7. OOD Experiments of LLMGuardrail
@@
Table 5 presents the results of out-of-domain (OOD) experiments for the LLMGuardrail model and several baseline methods on the ToxiGen dataset. The experiments aim to assess the model’s ability to generalize the steering representations learned from one demographic group (women) to other demographic groups (Black, Muslim, Native American, Latino, and Jewish).
@@
LLMGuardrail demonstrates strong generalization capability by effectively reducing toxicity and increasing refusal rates for all target demographic groups, even though the steering representations were learned using examples from the women group. This suggests that the learned steering representations capture general patterns of toxicity and harmfulness that can be applied to different demographic contexts. In addition, LLMGuardrail consistently outperforms the baseline methods (Base, LAT-Reading, LORRA, and Mean-Centring) across all target demographic groups, achieving higher refusal rates and lower toxicity percentages. This highlights the effectiveness of the causal analysis and adversarial learning techniques employed by LLMGuardrail in obtaining unbiased steering representations that generalize well to unseen demographic groups.
5.8. Sensitivity Analysis of LLMGuardrail
@@
Figure 6 and 7 present the sensitivity analysis of the hyperparameter $\alpha$ and $\beta$ on the ToxiGen and AdvBench datasets. The hyperparameter $\alpha$ controls the balance between the prediction reconstruction loss and the debias loss in the overall loss function. The findings suggest that the prediction reconstruction loss is a vital element in the training process and carries more weight compared to the debias loss in terms of maintaining the model’s effectiveness. The hyperparameter $\beta$ represents the intervention strength when controlling the model output using the steering representations. Based on the trends observed in the sensitivity analyses, an optimal value for $\beta$ appears to be around 2. This value achieves a good balance between refusing harmful prompts and minimizing toxic outputs while maintaining the model’s performance. The optimal ranges identified for $\alpha$ and $\beta$ can guide the selection of appropriate values for these hyperparameters in practical applications of the LLMGuardrail model.
5.9. Comparison with Aligned LLMs
Table 6 presents the experimental results of LLMGuardrail and aligned LLM (Wang et al., 2024a) on four benchmark datasets. We selected three aligned LLM methods, including supervised fine-tuning (SFT), and two RLAIF (Lee et al., 2023b) methods (Direct Preference Optimization, DPO(Rafailov et al., 2023) and Proximal Policy Optimization, PPO (Ouyang et al., 2022)). These alignment methods aim to enhance the model’s ability to generate safe and non-toxic content. Following Li et al. (2024), we constructed the SFT dataset and used GPT-4 to construct the preference dataset required for RLAIF. The experimental results demonstrate that our method outperforms SFT and DPO. Compared to PPO, our method shows greater advantages on Vicuna-7b and Llama2-7b, and it is worth noting that our algorithm does not require additional preference datasets and reward models, unlike PPO. Table 7 presents the results of LLMGuardrail and aligned LLM in OOD experiments, proving that LLMGuardrail can obtain unbiased steering representations with superior generalization performance than SFT, DPO, and PPO.
Table 8. Jailbreak attack experiments of our LLMGuardrail model on the AdvBench subset dataset.
LLM Method ASR w/o ASR w/ ASR
LLMGuardrail LLMGuardrail Reduce(%)
GCG 16.20 2.91 $\downarrow\textbf{82.03}$
AutoDAN 74.06 56.28 $\downarrow\textbf{24.01}$
Vicuna PAIR 30.15 20.26 $\downarrow\textbf{32.81}$
DeepInception 56.34 35.15 $\downarrow\textbf{37.60}$
5.10. Analyze the Effectiveness of LLMGuardrail on Jailbreak
We analyze the defense capability of LLMGuardrail against jailbreak attacks and utilize visualization methods to explain the effectiveness of LLMGuardrail. We selected four models for jailbreak attacks (GCG (Zou et al., 2023b), AutoDANp(Liu et al., 2023c), PAIR(Chao et al., 2023), and DeepInception(Li et al., 2023)), including three Optimization-based attacks and one Template-based attack(Wei et al., 2023). Following Li et al. (2023), we evaluate our methods on the subset in the AdvBench benchmark, the Vicuna-7B model as the base LLM, and employed Accack Success Rate (ASR) for evaluation. Table 8 presents the effective defense capabilities of LLMGuardrail against jailbreak attacks, reducing the ASR by an average of 44.11% across the four attack models.
We explain the effectiveness of the LLMGuardrail method within the representation space by examining the embedding distributions of different types of prompts in the hidden states of LLMGuardrail. Using t-SNE clustering, we visualize these distributions in Layer-7, as shown in Figure 8. This visualization illustrates the distribution of four types of prompts: those identified as safe by LLMs, those identified as unsafe by LLMs, representations of jailbreak attack prompts, and representations calibrated by LLMGuardrail.
Our analysis reveals a clear distinction between safe and unsafe prompts in the intermediate layer representation space of the LLM, demonstrating the model’s inherent ability to differentiate between these categories. However, a surprising and critical finding is the positioning of jailbreak attack prompts. These prompts are located near the boundary between safe and unsafe clusters, occupying a precarious position that allows them to potentially confuse the LLM. This strategic placement enables jailbreak attacks to bypass the model’s built-in safety mechanisms, potentially leading to the generation of harmful content. The boundary position of jailbreak attack prompts presents a significant vulnerability in LLMs. Due to their ambiguous location, these prompts may be mistakenly classified within the safe space. This misclassification can induce the LLM to incorrectly judge harmful prompts as harmless and respond to them accordingly, thereby compromising the model’s safety guardrails and producing potentially dangerous outputs. LLMGuardrail effectively shifts the representations of jailbreak attack prompts from their ambiguous boundary position back into the clearly defined unsafe space. This recalibration is visually represented by the yellow points in Figure 8, which show the adjusted positions of the jailbreak prompts after the application of LLMGuardrail. By repositioning these prompts, LLMGuardrail enables the model to correctly identify them as potentially harmful, preventing misjudgment and inappropriate responses.
The ability of LLMGuardrail to recalibrate prompt representations in the model’s embedding space is crucial for mitigating harmful behavior and improving the overall quality and safety of generated content. This recalibration allows the model to properly defend against or refuse to answer unsafe queries, thereby significantly enhancing its safety and reliability. Furthermore, this shift demonstrates LLMGuardrail’s capacity to adjust the model’s internal representations, making it more robust against inputs specifically designed to bypass safety measures.
6. Conclusion